OpenStudy (anonymous):

s^2-8s+c , Solve for c to make perfect square..

OpenStudy (freckles):

Recall: \[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2 \]

OpenStudy (freckles):

so what is your k here?

OpenStudy (anonymous):

Nvm, The answer to my question is 16.

OpenStudy (freckles):

correct! :)